Cavy Company estimates that total factory overhead costs will be $484,544 for the year. Direct labor hours are estimated to be 90,400.

a. Compute the predetermined factory overhead rate. Round your answer to the nearest cent.
$ per direct labor hour

b. Determine the amount of factory overhead applied to Job 345 if the amount of direct labor hours is 1,200 and to Job 777 if the amount of direct labor hours is 2,900.

Job 345 $
Job 777 $

c. Journalize the entry to record the factory overhead applied if Jobs 345 and 777 are the only jobs for the period. If an amount box does not require an entry, leave it blank.
Answer #1

Dear Student, Thank you for using homeworklib, please find below the answer and give a thumbs up.

Solution a) Computation of Predetermined O/H Rate
Predetermined O/H Rate = Estimated Factory O/H Costs
Estimated Direct Labor Hours
Predetermined O/H Rate = 5.4
(484544/90400)
Solution b) Computation of factory overhead applied to job 345 and job 777
Direct Labor Hours * Predetermined O/H Rate
Labour Hours Predetermined O/H Rate Applied O/H
Job 345 1200 5.4 6480
Job 777 2900 5.4 15660
Solution c) Journal entry to record to apply factory overhead:
Date Accounts and Explanation Debit Credit
Work in Progress 22140
              Factory O/H (6480+15660) 22140
(To record the material cost)